OpenStudy (anonymous):

how do i simplify roots ? here is an example: ( -√(7k) -6 )( 3√(7k) -6)

OpenStudy (anonymous):

just multiply the roots

OpenStudy (anonymous):

\[-\sqrt{7k}*3\sqrt{7k}=-3(7k)\]

OpenStudy (anonymous):

\[6\sqrt{7k}-18\sqrt{7k}=-12\sqrt{7k}\]

OpenStudy (anonymous):

\[6*6=36\]

OpenStudy (anonymous):

\[-21k-12\sqrt{7k}+36\]

OpenStudy (anonymous):

that is your solution did you understand it @eric.schult

OpenStudy (anonymous):

where did -21K come from ? m answer was: -15√(7k) +36

OpenStudy (anonymous):

the first step did you see it when you multiply roots the root sign goes away

OpenStudy (anonymous):

and as you can see -3*7 is 21

OpenStudy (anonymous):

i have never multiplied roots . i have only recently learned how to add and subtract them.

OpenStudy (anonymous):

the thing is when you multiply roots they cancel out

OpenStudy (anonymous):

so then for example : 5√(12) * 4√(12) becomes 20√(12) and then 20*12 ?

OpenStudy (anonymous):

where as : 3√(45) * 5 becomes 15√(45) and nothing more. correct ?

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

what about this situation? 4√(3x) * -7√(3) a variable exists in only 1

OpenStudy (anonymous):

now that is going to be \[-28\sqrt{3x}\sqrt{3}\]

OpenStudy (anonymous):

would the answer be : -28√(3^2x) and then -28*3√(x) ?

OpenStudy (anonymous):

excellent @eric.schult

OpenStudy (anonymous):

:) thanks again

OpenStudy (anonymous):

that is because \[\sqrt{3x}=\sqrt{3}\sqrt{x}\]